OpenStudy (anonymous):

Coefficient must be squared using the original x term and added to both sides of the equation 4x² - 8x = 52 solve

OpenStudy (anonymous):

Step 1: Put all ur terms on the same side. Therefore move 52 to the left side of the equation. what do you get when you do this?

OpenStudy (anonymous):

I would then add the 52 on the left side? Am I doing the opposite?

OpenStudy (anonymous):

Please explain... need your help.

OpenStudy (anonymous):

I am just walking you thru the steps. Can u do step 1?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

what do u get?

OpenStudy (anonymous):

If I move it to the left I get 52+4X2-8x=52

OpenStudy (anonymous):

or wait.... now I don't have terms on the right yet... ugh!

OpenStudy (anonymous):

when I say term on the right I mean the term that is to the right of the equal sign (52)

OpenStudy (anonymous):

Subtract 52 on both sides to move the 52 from the right of the equal (=) sign to the left side of the (=) sign. Step1: 4x^2-8x=52 -52 -52 You should get: 4x^2-8x-52=0

OpenStudy (anonymous):

Do u understand this?

OpenStudy (anonymous):

I need you to explain why?

OpenStudy (anonymous):

We are trying to solve for 'x' In order to do that, we need to move the 52 to the left hand side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is there a common factor between the coefficients (or numbers) in the quadratic equation?

OpenStudy (anonymous):

huh? The gcm?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

I'm dumb at this stuff

OpenStudy (anonymous):

Close nope its 4. So take 4 outside a parentheses: 4(x^2-2x-13)=0

OpenStudy (anonymous):

Divide both sides by 4: 4/4(x^2-2x-13) = 0/4 Get: x^2-2x-13 = 0 next step is to factor: by factoring or use the quadratic formula.

OpenStudy (anonymous):

Do u know the quadratic formula and how to use it to solve a quadratic equation?

OpenStudy (anonymous):

Quadratic formula: \[x=-b \pm \frac{ \sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

Correct equation: \[x=\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\] a=1, b= -2, c=-13 \[x=\frac{-(-2)\sqrt{(-2)^{2}-4(1)(-13)} }{ 2(1) }\]

OpenStudy (anonymous):

Do you understand how i got this? 1x^2-2x-13=0, a=1, b=-2, c=-13, note got it from the numbers without the x-variable

OpenStudy (anonymous):

r u there? @asrodgers